Please help me solve this linear system:
4x-y=-1
6x+5y=18

I need to find out what would be x and y..
Thanks

Answer by drk(1908)

Here are your two original equations
(i) 4x - y = -1
(ii) 6x + 5y = 18.
There are several ways to solve. but I will use elimination.
multiply (i) by 5 to get
(iii) 20x - 5y = -5
(ii) 6x + 5y + 18
add down to get rid of the y's and solve for x.
(iv) 26x = 13
x = 1/2.
find y by placing 1/2 for x into (i) to get
4(1/2) - y = -1
2 - y = -1
y = 3.
Your solution is (1/2,3)